n=36 x=24.6 s=12 H0: m<=20 Ha: m>20 If the test is done at a .05 level...

n=36

x=24.6

s=12

H0: m<=20

Ha: m>20

If the test is done at a .05 level of significance, the null hypothesis should _____.

a. not be rejected

b. be rejected

c. Not enough information is given to answer this question.

d. None of these answers are correct.
Answer #1

SE = s/sqrt{n}

= 12/V36 = 2

Test statistic is:

t = (24.6 - 20)/2 = 2.3

alpha =0.05

ndf = 36- 1 = 35

One Tail - Right Side Test

From Table, critical value of t = 1.6896

Since the calculated value of t = 2.3 is greater than critical value of t = 1.6896, the difference is significant. Reject null hypothesis.

So,

Correct option:

b. be rejected.